OpenStudy (anonymous):

What is the slope of a line that passes through the point (−1, 1) and is parallel to a line that passes through (3, 6) and (1, −2)?

OpenStudy (anonymous):

lines that are parallel have the same slope, so first you want to find the slope of the line that passes through (3,6) and (1,-2)

OpenStudy (anonymous):

you know how to find slope?

OpenStudy (anonymous):

y1-y2 over x1-x2 or something right? @juliannescriv

OpenStudy (anonymous):

yep :) either \[\frac{ y _{1}-y _{2} }{ x _{1}-x _{2}} \] or have the 2s on the left and the ones on the right

OpenStudy (anonymous):

so do \[\frac{ 6-(-2) }{ 3-1 }\]and tell me what you get

OpenStudy (anonymous):

8/2

OpenStudy (anonymous):

and that simplifies to 4

OpenStudy (anonymous):

point slope is\[y-y _{1}=m(x-x _{1})\] \[y _{1}=6\]\[x _{1}=3\]

OpenStudy (anonymous):

plug those numbers in and solve and tell me what you get

OpenStudy (anonymous):

oh, and m=4

OpenStudy (anonymous):

you understand where i got 6, 3, and 4?

OpenStudy (anonymous):

yes, where u found the slope. so y-6=4(x-3) ?

OpenStudy (anonymous):

yep. next you want to distribute that 4

OpenStudy (anonymous):

y-6=4x-12

OpenStudy (anonymous):

great. last you want to get y by itself

OpenStudy (anonymous):

you know how to do that?

OpenStudy (anonymous):

add 6 so y=4x-6 ?

OpenStudy (anonymous):

ahh, so sorry. i messed up

OpenStudy (anonymous):

we were supposed to use (−1, 1) instead of (3,6)

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

so, use that point slope equation again, y= 1, x=-1, m=4

OpenStudy (anonymous):

but you got it the first time, im sure you can do it again :P

OpenStudy (anonymous):

y-1=4(x+1)?

OpenStudy (anonymous):

distribute the 4 again

OpenStudy (anonymous):

y-y=4x+4

OpenStudy (anonymous):

*y-1

OpenStudy (anonymous):

next get y by itself

OpenStudy (anonymous):

ya, y-1 sorry. so y=4x+3

OpenStudy (anonymous):

nope, you add the 1

OpenStudy (anonymous):

ohh my bad 5

OpenStudy (anonymous):

good job

OpenStudy (anonymous):

so what do i do now?

OpenStudy (anonymous):

m=4

OpenStudy (anonymous):

wasn't sure if you needed the equation for that line too or not

OpenStudy (anonymous):

ohh okay! thanks!!!!!

OpenStudy (anonymous):

mhm :)

OpenStudy (anonymous):

how do i give medals on here lol

OpenStudy (anonymous):

click best response :P